### v2

# Octopus Inspired Optimization Algorithm: Multi-Level Structures and Parallel Computing Strategies

Abstract
This paper introduces a novel bionic intelligent optimisation algorithm, Octopus Inspired Optimization (OIO) algorithm, which is inspired by the neural structure of octopus, especially its hierarchical and decentralised interaction properties. By simulating the sensory, decision-making, and executive abilities of octopuses, the OIO algorithm adopts a multi-level hierarchical strategy, including tentacles, suckers, individuals and groups, to achieve an effective combination of global and local search. This hierarchical design not only enhances the flexibility and efficiency of the algorithm, but also significantly improves its search efficiency and adaptability. In performance evaluations, including comparisons with existing mainstream intelligent optimisation algorithms, OIO shows faster convergence and higher accuracy, especially when dealing with multimodal functions and high-dimensional optimisation problems. This advantage is even more pronounced as the required minimum accuracy is higher, with the OIO algorithm showing an average speedup of 2.27 times that of conventional particle swarm optimisation (PSO) and 9.63 times that of differential evolution (DE) on multimodal functions. In particular, when dealing with high-dimensional optimisation problems, OIO achieves an average speed of 10.39 times that of DE, demonstrating its superior computational efficiency. In addition, the OIO algorithm also shows a reduction of about 5% in CPU usage efficiency compared to PSO, which is reflected in the efficiency of CPU resource usage also shows its efficiency. These features make the OIO algorithm show great potential in complex optimisation problems, and it is especially suitable for application scenarios that require fast, efficient and robust optimisation methods, such as robot path planning, supply chain management optimisation, and energy system management.
Keywords Heuristic Algorithms; Bionic Optimisation; Octopus-inspired algorithms (OIO); Distributed processing; Asynchronous Parallel Mechanisms
1. Introduction
Heuristic algorithms are a class of optimization algorithms that mimic biological or physical phenomena found in nature [1-4]. These algorithms seek approximate optimal solutions to complex problems by emulating mechanisms such as evolution, group behaviors, and immune system functions. Compared to traditional mathematical planning methods, heuristic algorithms can often find satisfactory solutions in a shorter time and effectively avoid the problem of falling into local optima.
Among these, Swarm Intelligence (SI) is one of the most well-known branches of heuristic algorithms [7-8]. SI simulates the collective behaviors of organisms in nature, such as ants, flocks of birds, and schools of fish, achieving group optimization through simple interactions between individuals. Classic SI algorithms include Ant Colony Optimization (ACO) [9-10], Particle Swarm Optimization (PSO) [11-12], and others. These algorithms showcase the great potential of heuristic approaches in solving complex optimization problems. However, existing bionic intelligence optimization algorithms primarily focus on the "emergence" of intelligence from cooperative behaviors between individuals, with less emphasis on the role of individual intelligence in this process. In fact, the intelligent behaviors of many organisms are shaped by intricate neural structures. For instance, ants have numerous olfactory receptors on their antennae, and the ganglia in their abdomen integrate multiple sensory inputs [13-14]. These complex perceptual and neural computational mechanisms form the basis of ants’ intelligent behavior. Therefore, we believe it is essential to consider the neuromorphological characteristics of individuals more thoroughly when designing bionic intelligence optimization algorithms.
In this paper, we propose a new bionic intelligence optimization algorithm inspired by the unique behavioral features of the octopus, such as foraging and predator avoidance, and the hierarchical structure of the octopus nervous system. The algorithm is designed around key features of the octopus, including the hierarchical structure of the suckers, the decentralized interaction of the tentacles, and information exchange between individuals. Mechanisms such as multi-level search, asynchronous parallelism, adaptive parameter control, and region regeneration are incorporated. Through reasonable abstraction and mapping, we construct a flexible and efficient Octopus Inspired Optimization (OIO). Compared to traditional swarm intelligence algorithms, this work focuses more on extracting intelligent mechanisms from neuromorphological features, aiming to enhance the adaptability and emergent abilities of the algorithm. This approach can avoid local optima while maintaining algorithmic simplicity, demonstrating excellent performance in various benchmark functions and practical application problems.
2. Related Work
In recent years, many researchers have proposed various novel optimization algorithms inspired by biological mechanisms in nature. For example, the Particle Swarm Optimization (PSO) algorithm proposed by Kennedy and Eberhart in 1995 simulates the flocking behavior of birds when foraging for food and achieves optimization through information exchange between individuals. Dorigo et al. drew on the pheromone mechanism of ants searching for food and proposed the Ant Colony Optimization (ACO) algorithm in 2006. Karaboga proposed the Artificial Bee Colony (ABC) algorithm in 2005, inspired by the honey harvesting behavior of honey bees. The ABC algorithm cleverly balances the relationship between exploration and exploitation, showing good optimality searching ability on several benchmark functions[15-16]. Mirjalili et al. proposed the Grey Wolf Optimizer (GWO) in 2014, inspired by the grey wolfs hunting mechanism. GWO demonstrated excellent performance in solving multiple benchmark functions and engineering problems by simulating the social hierarchy and hunting behavior of grey wolf packs [17-18]. Xu et al. proposed the Bat Sonar Algorithm (BSA) in 2021 [19-20]. Although these algorithms show the superiority of group intelligence, they mainly focus on modeling individual behaviors and lack sufficient exploration of the organizational structure of the group. For example, the interaction between particles in the PSO algorithm is relatively simple, making it difficult to portray the complex group decision-making process [21-22].
Although the above algorithms have achieved notable results in their respective fields, they either focus only on modeling the behavior of individual organisms and lack in-depth exploration of the higher intelligence of the group, or they qualitatively focus on individual organisms without adequately addressing the structural complexity of group relationships. In contrast, octopuses possess a unique hierarchical and autonomous nervous system, including multi-layered structures such as suckers, tentacles, and individuals, which support their flexible and varied sensing, decision-making, and execution abilities. Additionally, there is a certain degree of synergistic cooperation and optimization among groups [23-24]. Inspired by this, this paper proposes the Octopus Inspired Optimization (OIO) to achieve a breakthrough in intelligent optimization algorithms by simultaneously considering individual learning and group cooperation.
The innovations of the Octopus Inspired Optimization (OIO) include:
1. Proposal of a New Bionic Intelligence Optimization Algorithm: Inspired by the nervous system of the octopus, the OIO algorithm designs mechanisms such as multi-level search, asynchronous parallelism, and adaptive parameter control to fully utilize the octopus’s perception, decision-making, and execution capabilities.
2. Algorithm Implementation Strategy: Based on various neuromorphological features of the octopus, such as suckers, tentacles, individuals, and swarms, specific implementation strategies for the OIO algorithm are tailored.
3. Performance Evaluation: The OIO algorithm is compared with existing mainstream intelligent optimization algorithms in a multi-dimensional benchmark test. The results show that the OIO algorithm excels in convergence speed, solution quality, and robustness.
4. Scalability and Application Prospects: The bionic design of the OIO algorithm can be generalized to other intelligent computing models such as neural network design and has potential applications in various fields such as pattern recognition, robot control, and intelligent optimization.
Subsequent chapters of this paper will detail the design of the OIO algorithm. Chapter 3 will explore the morphological features of the octopus nervous system and their implications for algorithm design. Chapter 4 will detail the framework and methodological principles of the OIO algorithm, focusing on the specific implementations of multi-level search, asynchronous parallelism, and adaptive control. Chapter 5 will systematically evaluate the performance of the OIO algorithm through simulation experiments and analyze it in comparison with other intelligent optimization algorithms.
3. Morphological Feature Analysis of the Octopus
Octopuses are highly evolved cephalopod mollusks with complex nervous systems and remarkable intelligence. Their nervous system comprises approximately 500 million neurons, most of which are distributed in their tentacles, enabling them to perform functions such as grasping and exploration independently [25-27]. The suckers on octopus tentacles are embedded with multimodal receptors that can integrate and transmit environmental information to higher neural centers [28]. These features allow octopuses to interact precisely with their environment, providing important insights for the design of intelligent algorithms.
Overall, the information processing flow of the octopus nervous system is as follows: the suckers selectively perceive environmental information, which is integrated and summarized in the ganglia of the tentacles. Based on proprioceptive information and inputs from other tentacles, the tentacles independently execute various behaviors such as movement, extension, growth, and retraction. The individual octopus gains global information and controls and adjusts the movement of the tentacle structure accordingly. At the group level, when information is obtained from different octopuses, it is possible to abstract an overall profile of the problem solution. Based on the needs of problem-solving, the group adjusts through master-slave co-evolution to obtain targeted information.
3.1 Octopus Basic Nervous System
The nervous system of the octopus has a unique structure, from the suckers to the ganglia and then to the tentacles, as shown in Figure 1. This paper systematically analyzes the key features at these structural levels and their impact on algorithm design.
Figure 1: Diagram of the Nervous System of Octopuses
Inspired by the morphological characteristics of the octopus, we abstract four levels: sucker, tentacle, individual, and group to simulate the emergence of intelligent behavior. After aggregating individual information, different information sources from octopuses are integrated to establish an abstract concept of the group, achieving multimodal source integration.
3.2 Suckers: Information Perception and Local Processing
In group optimization problems, we use multidimensional function values to evaluate the function of the sucker structure, simulating the independent control of suckers by the tentacles. This design allows suckers to perform efficient local information perception and processing, providing necessary feedback to the upper structures, thereby demonstrating the distributed processing characteristics of the octopus nervous system. The specific details of information transmission and parameter adjustment are shown in Figure 2.
In the algorithm design, the suckers are initialized at the beginning of each iteration. They first collect information based on the evaluation function, then aggregate and feedback this information to the tentacle structure. Before the next iteration, the tentacle structure adjusts the relevant parameters, updates the performance indicators of the suckers, and optimizes the information processing process, outputting the aggregated information to the upper structures. This process can be represented by the following formula:
${S}_{i}^{t+1}=f\left({{S}_{i}^{t},{E}_{i}^{t},{P}_{i}^{t}}\right)$ (1)
where ${S}_{i}^{t}$ represents the state of the $i$ th sucker at iteration $t,{E}_{i}^{t}$ represents the environmental information it acquires, ${P}_{i}^{t}$ represents the control parameters returned by the tentacle structure, and $f$ is the information processing function of the sucker.
Figure 2: Diagram of the Suckers
3.3 Tentacles: Distributed Search and Autonomous Coordination
During each iteration, the tentacles first integrate information from the suckers and, after evaluating the environment, make autonomous decisions involving movement adjustments (e.g., moving, extending) and sucker function regulation. All information is ultimately fed back to the individual, influencing the motion instructions for the next iteration. This process can be represented by the following formula:
${T}_{i}^{t+1}=g\left({{T}_{i}^{t},{S}_{i}^{t},{I}^{t}}\right)$ (2)
where ${T}_{i}^{t}$ is the state of the $i$ th tentacle, ${S}_{i}^{t}$ is the set of sucker states, ${I}^{t}$ is the motion.
Figure 3: Diagram of the Tentacles
As shown in Figure 3. In group optimization, the behavior of the tentacles is simulated by internal sucker particles, adjusting the search range and position of the suckers on the tentacles to simulate tentacle extension and retraction. The tentacles elongate when moving towards optimal regions and contract or degenerate otherwise, balancing the exploration and exploitation of the algorithm to optimize performance. The growth and degeneration mechanisms of the tentacles need to be designed according to the problem and parameters, selecting appropriate trigger conditions and adjustment strategies to ensure the stability and efficiency of the algorithm while considering computational overhead to avoid performance degradation.
3.4 Individuals: Macro Decision-Making and Adaptive Regulation
In the Octopus Inspired Optimization, individuals not only act as macro decision-makers and adaptive regulators by integrating tentacle information and making decisions based on environmental characteristics, but also control the group optimization algorithm through a reward and punishment mechanism. Specifically, individuals reward or punish tentacles based on their performance in the global environment, and reinitialize tentacles that remain in low-value or local optimal regions for extended periods. This control structure allows individuals to guide the group optimization algorithm to adaptively adjust search strategies, thereby enhancing overall algorithm performance. This process can be represented by the following formula:
${I}^{t+1}=h\left({{I}^{t},{T}^{t},{M}^{t}}\right)$ (3)
Figure 4: Diagram of the Octopus Individuals
Where ${I}^{t}$ is the individual state, ${T}^{t}$ is the set of tentacle states, ${M}^{t}$ is the memory information, and $h$ is the decision function. By dynamically adjusting the forgetting factor $\epsilon$ and the memory factor $\gamma$, individuals use the update function $\phi$ to balance the preservation of historical information with the forgetting of past optimal values, optimizing search efficiency and algorithm convergence. The process of individuals adaptively adjusting search strategies is illustrated in Figure 4.
3.5 Group Intelligence: Emergent Properties and Collaborative Optimization
The group aggregates information from octopus individuals of the same or different "types" and synthesizes multiple data sources to comprehensively describe the problem-solving process. For each octopus’s data, the group abstracts the information and integrates all types of data into the same dimension. For each octopus’s movement, the group regulates based on the needs of problem-solving, enabling each octopus to collaboratively address the issue.
In the group optimization algorithm, since the input data are merely the extreme value solutions of multidimensional functions, we establish a reinforcement learning model and a global optimal control mechanism for each octopus. This design allows the algorithm to gain stronger regulation capabilities (e.g., avoiding local optima) and more comprehensive search abilities for the global optimum during the search process.
4. Methodology of the Octopus Inspired Optimization
4.1 Group Co-evolution
The Octopus Intelligence Optimization (OIO) algorithm is inspired by the high cooperation capability of octopus groups, adopting a master-slave group co-evolution strategy. This strategy decomposes the overall problem into multiple sub-problems, which are solved through information sharing and cooperation among sub-groups, thereby overcoming the local optimal traps commonly encountered by traditional algorithms in high-dimensional complex problems[29-30].
Figure 5: Diagram of Group Co-evolution
(A) Master points searching for areas of dominance;
(B) Some subject points finding areas of dominance and guiding other follower points to adjust their stride and direction;
(C) Search points close to the search area switching from follower object to subject object.
As shown in Figure 5, the main group deeply exploits the current optimal region, while the subordinate group extensively explores new areas. When the subordinate group discovers a potential area, it converts into the main group through a "guidance factor," and the main group adjusts the search direction based on the concentration of the "guidance factor." The algorithm also includes a master-slave identity switching mechanism to maintain a balance between exploration and exploitation [4]. The pseudocode for the algorithm’s partial principles is shown in Pseudocode 5.
4.2 Memory-Based Evolution Strategy
The memory and learning abilities of octopuses in nature also provide inspiration for the design of the OIO algorithm. Our memory-based strategy uses evolutionary historical information such as individual heredity and fitness statistics to guide the evolution process, balancing global and local searches [31-32]. The algorithm maintains a population memory bank that records optimal solutions and decision paths, allowing octopus individuals to use the memory bank to guide their search direction. To prevent excessive reliance on specific strategies, a tentacle forgetting mechanism is introduced:
${M}_{i}^{t+1}=\lambda{M}_{i}^{t}+\left({1-\lambda}\right){\bar{M}}_{i}^{t}$ (4)
Where ${M}_{i}^{t}$ is the tentacle memory, ${\bar{M}}_{i}^{t}$ is the new memory, and $\lambda$ is the forgetting coefficient that controls memory retention and forgetting, maintaining strategy diversity. Additionally, we integrate the forgetting mechanism with the tentacle growth and degradation mechanism. When dependence exceeds a threshold, the tentacle degrades, promoting the growth of new tentacles, and maintaining global search capability [33-34]. This design enhances the algorithm’s robustness and global optimization capability.
By designing an effective tentacle forgetting mechanism combined with the growth and degradation mechanism, the OIO algorithm retains useful prior knowledge while avoiding overreliance on a single strategy, thereby improving the algorithm’s robustness and global optimization capability. The core code is shown in Pseudocode 2.
The Octopus Inspired Optimization (OIO) algorithm, inspired by the octopus nervous system, designs a localized structure simulating the division of labor between suckers and tentacles, improving global optimization efficiency.
In the OIO algorithm, suckers perform local search and direction guidance, with each sucker representing a point in the solution space partition, focusing on local sampling and exploitation.
The workflow includes: initializing with randomly deployed suckers, each sucker performing local perturbations and position updates based on location and objective function values, and the updated suckers interacting through a tentacle mechanism to guide the global search direction.
The tentacle structure is abstracted into a spatial plane covering parts of the solution space, responsible for overall search and information transmission. Multiple tentacles are generated during initialization to cover the solution space. Tentacles adjust their shape through interactions with suckers, exchanging information and influencing each other to achieve efficient overall search.
As shown in Figure 6, the tentacles work in tandem with the suction cups and will dynamically adjust the search area of the tentacles during the iteration process.
Figure 6: Schematic diagram of the iterative process
Through the collaboration of suckers and tentacles, the OIO algorithm maintains a dynamic balance between fine-tuning known good solutions (exploitation) and broadly exploring unknown areas (exploration), enhancing performance. This coordination mechanism fully utilizes the octopus nervous system’s high division of labor and precise control characteristics, providing a novel bio-inspired solution for complex optimization problems. The core process of the algorithm can be represented by the Pseudocode 3.
4.4 Adaptive Search and Diversification Strategy
The Octopus Inspired Optimization (OIO) algorithm integrates adaptive search mechanisms and diversification strategies to enhance its adaptability to complex environments and avoid premature convergence.
daptive Search Mechanism: Includes tentacle regeneration and plasticity strategies. Tentacle regeneration is achieved through the condition $f\left({x}_{i}\right)<{\theta}_{r}$, allowing tentacles to regenerate in promising areas and reduce ineffective searches. The plasticity strategy adjusts weights through:
$w_{j}=score_{j}/\sum_{k=1}^{N_{s}}score_{k}$ (5)
enabling the algorithm to autonomously adjust strategy combinations based on previous performance.
Diversification Strategy: includes asynchronous iteration queue search and multi-objective optimization guidance. Asynchronous queue search mode:
${x}_{i}^{t+1}=\left\{\begin{matrix}{x}_{i}^{t}&{\text{if }i\notin{Q}^{t}}\\ {\mathcal{U}({x}_{i}^{t})}&{\text{if }i\in{Q}^{t}}\end{matrix}\right.$ (6)
enhances population diversity, avoiding concentration on non-global optima due to synchronous updates. Multi-objective optimization:
$\mathop{\min}\limits_{{x\in\Omega}}\mathbf{F}\left(x\right)={\left({f}_{1}\left(x\right),{f}_{2}\left(x\right),\ldots,{f}_{m}\left(x\right)\right)}^{\mathrm{T}}$ (7)
helps the algorithm maintain population diversity while effectively avoiding excessive convergence to boundaries.
Through adaptive adjustments of tentacle regeneration and plasticity strategies, combined with asynchronous iteration and multi-objective optimization, the OIO algorithm demonstrates excellent global search capabilities and robust convergence performance, effectively addressing dynamic and complex optimization problems.
5. Experimental Section
5.1 Basic Information of the Experiment
To evaluate the performance of the Octopus Inspired Optimization (OIO) algorithm, we selected a series of representative test functions, such as the multimodal Rastrigin function and the unimodal Easom function. These test functions are diverse and representative, commonly used in the industry to examine the performance of algorithms in handling different types of optimization problems. As shown in Table 1 below, we recorded their mathematical expressions, function names, and function domains according to their classifications. Based on the dimensional characteristics of the functions, we categorized the test functions into four types: Fixed Dimension (F), Random Dimension(R), Unimodal (U), and Multimodal (M).
Fixed dimensionality functions are used to test the algorithm’s ability to optimise at a specific complexity; random dimensionality functions assess the algorithm’s adaptability and flexibility with respect to dimensional changes; unimodal functions help analyse the algorithm’s efficiency in local search, while multimodal functions are used to verify the algorithm’s ability to avoid local optimums and to explore the potential solution space in a global search. This categorisation ensures that the performance of OIO algorithms in all standard scenarios is fully and fairly tested.
Additionally, to better compare the time efficiency, resource utilization, and other parameters of the algorithms horizontally, we also selected several classic heuristic algorithm models such as Particle Swarm Optimization (PSO) and Genetic Algorithm (GA) as control groups. The control groups were tested multiple times in parallel under the same server environment and with the same hyperparameters to ensure fairness. The details of the operating system and hardware configuration are as follows:
Operating System and Software Environment:
  • Python 3.7
  • CUDA 11.1
  • Pytorch 1.8.1
  • Ubuntu 18.04
  • NVCC 11.1
  • OpenMPI 4.0.0
Speed Test Hardware Configuration:
  • GPU: NVIDIA GeForce RTX 3090, 24GB Graphics Memory
  • CPU: Intel Xeon Gold 6230R Processor, 24 cores, 35.75M Cache, 2.10 GHz
  • Memory: 120GB RAM
  • Storage: 1500GB ROM
Accuracy Test Hardware Configuration:
  • CPU: Intel Xeon Gold 6230R Processor, 1 core, 35.75M Cache, 2.10 GHz
  • Memory: 4GB RAM
  • Storage: 150GB ROM
Table 1: Test Functions
[…]
5.2 Experimental Parameters and Performance Testing
Considering the requirements of actual production and physical realities, our speed and accuracy tests were conducted in a three-dimensional physical field to solve extremum problems.
5.2.1 Speed Test Setup
To ensure efficiency while controlling the minimum search unit number comparable to other algorithms, we set the following experimental parameters:
Number of Iterations: 50 iterations for the population, a number sufficient for convergence across various algorithms.
Octopus Configuration: Initialization with 4 octopuses, each equipped with 5 tentacles.
Tentacles and Suckers: Each tentacle is equipped with 50 suckers, and each tentacle undergoes 200 iterations.
Adaptive Parameter Adjustment: Given that the OIO algorithm employs adaptive hyperparameter adjustment, we set thresholds for related hyperparameters (total_range $=$ max_range - min_range, where max_range and min_range are the upper and lower limits of the search space, respectively). The hyperparameter ranges are shown in Table 2:
Table 2: Parameter Constraints
Tentacle Rebirth Parameter (20,40,0.06,0.08) Expansion Factor Range (0.4,1.6) Radius Range $(0.02\times\text{total}_{\text{range}},0.3\times\text{total}_{\text{range}})$ Number of Individual Tentacle Suckers (10,400) Range of Tentacle Iterations (20,400)
To test the speed performance of the OIO algorithm, we set acceptable threshold values for the search results at $1\mathrm{e}-{10},1\mathrm{e}-{12},1\mathrm{e}-{14}$ , and $1\mathrm{e}-{16}$ The search ends as soon as the algorithm finds a point meeting the threshold condition, and the time consumed for this process is recorded. These thresholds represent 5 progressive experimental groups, each group undergoing 10 parallel tests consecutively.
5.2.2 Accuracy Test
In the accuracy test, to increase the experimental difficulty and demonstrate the performance of a single octopus in a two-dimensional search problem, we reduced the configuration from the original four octopuses to a single octopus. This adjustment was made to investigate the robustness and effectiveness of the OIO algorithm under more challenging conditions. The experiment finely adjusts the number of tentacle iterations and suckers, dividing it into the following three experimental groups, with each group undergoing ten independent tests to ensure the reliability and repeatability of the results:
(a) Experimental Group 1: Single tentacle with 50 suckers, 200 iterations.
(b) Experimental Group 2: Single tentacle with 40 suckers, 180 iterations.
(c) Experimental Group 3: Single tentacle with 30 suckers, 160 iterations.
Under these settings, we adjust the step size for tentacle iterations and the number of suckers to 20 and 10, respectively, to accurately evaluate the algorithm’s sensitivity to parameter changes. This testing configuration allows us to thoroughly investigate the algorithm’s optimization capabilities under various parameter settings, particularly its performance under extreme or boundary conditions.
5.3 Speed Benchmark and Methodology
This study involves various population optimization algorithms, including Particle Swarm Optimization (PSO), Artificial Bee Colony (ABC), Genetic Algorithm (GA), Artificial Immune System (AIS), Differential Evolution (DE), and Octopus Inspired Optimization (OIO). To ensure consistency and fairness, the total number of particles for all algorithms is uniformly set to 1000 (calculated as 4*5*50), and the total number of iterations is set to 10000 (calculated as 50*200). Referencing the research by D. Karaboga et al. [1], in this experiment, the mutation rate for GA is set to 0.01, the crossover rate to 0.9, and the generational gap selection to 0.9 . For PSO, the cognitive and social influence factors are both set to 1.8, and the inertia weight is set to 0.6 .
The acceptable threshold values for speed testing are set to 1e-8, 1e-10, 1e-12, 1e-14, and 1e- 16. Whenever an algorithm finds a solution below these thresholds, the search is immediately terminated and the time consumed is recorded. For each threshold, all algorithms undergo 10 independent continuous runs to ensure the reliability and consistency of the results.
5.4 CPU Usage Test
Given that AIS, ABC, and GA did not find solutions meeting the threshold in previous experiments, this part of the test will exclude these three algorithms. Although DE was able to find solutions meeting the threshold, it took too long, rendering it unsuitable for further CPU usage comparison. This section aims to investigate whether the speed improvement of the OIO algorithm is due to higher CPU usage. To this end, we designed an experiment where both OIO and PSO algorithms optimize 12 test functions under the set acceptable threshold of 1e-14, recording CPU usage throughout the experiment. This experiment not only measures algorithm performance but also evaluates their resource efficiency in practical operations, providing a more comprehensive perspective on the applicability of the algorithms in real-world scenarios.
5.5 Engineering Problems
In this section, we use the OIO algorithm to solve two typical engineering design problems: three-bar truss design [35] and tension/compression spring design [36]. In order to ensure the reliability of the experimental results and the stability of the algorithm performance, we set the population size of the OIO algorithm to 30 to ensure the diversity and control the computational cost, the number of iterations was set to 25 to balance the convergence speed and optimisation effect, and 20 independent runs were conducted to take the average value in order to reduce the influence of random factors and enhance the statistical significance of the results.
5.5.1 The Three-member Truss Design Problem
The three-member truss design problem is a classic nonlinear fractional optimization issue in the field of civil engineering. The objective is to minimize the volume of the truss while ensuring its mechanical performance, thereby achieving an economically efficient use of materials. To accomplish this, various constraints must be imposed on each truss member, including buckling constraints, deflection constraints, and stress constraints. The schematic Figure 7 below illustrates the three-member truss problem, which involves two key parameters: cross-sectional areas $A_{1}$ and $A_{2}$.
Figure 7: Schematic Diagram of the Three-Member Truss Problem
a) Design Variables
$\mathbf{x}=\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}=\begin{bmatrix}A_{1}\\ A_{2}\end{bmatrix},\quad 0\leq x_{1},x_{2}\leq 10$ (2)
Here, $x_{1}$ and $x_{2}$ represent the cross-sectional areas $A_{1}$ and $A_{2}$ of the truss members, respectively.
b) Objective Function The goal is to minimize the total volume of the truss, expressed as:
$\text{Minimize}\quad f(\mathbf{x})=\left(2\sqrt{2}\,x_{1}+x_{2}\right)\cdot l_{0}$ (3)
where the constant $l_{0}=100\,\text{cm}$ denotes the length of the truss.
c) Constraint Conditions To ensure the structural safety and performance of the truss, the following three constraints must be satisfied:
[…]
In these equations, the constants $P=2\,\text{km/cm}^{2}$ represents the applied load, and $\sigma=2\,\text{km/cm}^{2}$ is the stress limit.
5.5.2 The Tension/Compression Spring Design Problem
The tension/compression spring design problem is a classic structural engineering optimization challenge. The objective is to minimize the weight of the spring while satisfying four constraints related to deflection, shear stress, vibration frequency, and outer diameter. To address this problem, three key variables are considered: wire diameter ($d$), mean coil diameter ($D$), and the number of active coils ($N$). The specifics of the spring and these three parameters are illustrated in Figure 8 below. The mathematical model of this problem is defined as follows:
Figure 8: Schematic Diagram of the Tension/Compression Spring Problem
a) Design Variables
$\mathbf{x}=\begin{bmatrix}x_{1}\\ x_{2}\\ x_{3}\end{bmatrix}=\begin{bmatrix}d\\ D\\ N\end{bmatrix},\quad 0.05\leq x_{1}\leq 2,\quad 0.25\leq x_{2}\leq 1.30,\quad 2\leq x_{3}\leq 15$ (5)
Here, $x_{1}$, $x_{2}$, and $x_{3}$ represent the wire diameter ($d$), mean coil diameter ($D$), and the number of active coils ($N$), respectively.
b) Objective Function The goal is to minimize the weight of the spring, expressed as:
$\text{Minimize}\quad f(\mathbf{x})=(x_{3}+2)\cdot x_{2}\cdot x_{1}^{2}$ (6)
This function aims to minimize the weight of the spring based on the design variables.
c) Constraint Conditions To ensure the structural safety and performance of the spring, the following four constraints must be satisfied:
[…]
  • Constraint $g_{1}(\mathbf{x})$: Ensures that the deflection does not exceed the allowable limit.
  • Constraint $g_{2}(\mathbf{x})$: Relates to the shear stress within the spring to prevent failure.
  • Constraint $g_{3}(\mathbf{x})$: Maintains the vibration frequency within desired bounds.
  • Constraint $g_{4}(\mathbf{x})$: Limits the outer diameter of the spring to fit within specified dimensions.
6. Results
6.1 Octopus Inspired Optimization Performance
The OIO speed test is measured in terms of elapsed time and its mean and standard deviation are shown in Table 3.
When the acceptable threshold is decreasing, the OIO elapsed time shows a more significant increase in some of the functions (e.g., Rosenbrock, Trid, Griewank, etc.), but all of them are able to find the acceptable threshold.
The OIO accuracy test was performed 10 consecutive times for each function separately, and the statistical means and standard deviations are shown in Table 4. (50,200) indicates a tentacle with 50 suckers and 200 iterations; the same is true for the remaining parameters.
According to Table 4 it can be found that the OIO algorithm has excellent performance in terms of accuracy for all types of test functions in 2 dimensions and is extremely stable.
The convergence plot of the octopus optimisation algorithm for different test functions in the accuracy test is shown in the second column of figure 6 . To ensure that the OIO algorithm does not fall into any other local minima in the accuracy experiments, we visualise the positions of the final suction cups and tentacles. x represents the position of the centre of the tentacle, the square box represents the tentacle coverage, the red dots represent the suction cups, and the star-shaped indicator represents the target position. An example of $\left({{50},{200}}\right)$ is shown in the third column of Figure 7.
Figure 9: Search Result
Figure 9 shows the convergence plot of the octopus optimisation algorithm for different test functions in the accuracy test. Based on the visualisation it can be found that the OIO algorithm does not overlap the tentacle positions at the same point when running. This is due to the region regeneration competition mechanism designed in the OIO algorithm. Although there is still some overlap in the region covered by the tentacles, the tentacles with less adaptation will be forced to randomly respawn at any point in the search space in the next iteration. This not only avoids the situation of multiple tentacles searching the same region repeatedly, but also improves the ability of the OIO algorithm to jump out of local minima.
6.2 Speed cross-comparison results
Tables 5, 6, 7, 8 and 9 summarise the mean and standard deviation of the elapsed time for each algorithm for each test function and compare the elapsed time while ensuring that acceptable thresholds were found, and the data with the lowest elapsed time will be marked in bold. Referring to the experimental tables in the appendix it can be seen that GA, AIS, and ABC do not satisfy the thresholds in the vast majority of cases and therefore are not involved in the comparison.
Figure 10: Search Speed Comparison Graph
Figure 10 shows the speed comparison graphs obtained in the four sets of speed tests. We can intuitively see that in the vast majority of cases, OIO takes less time than PSO and DE. in the four sets of tests, OIO achieves an average speed of 1.69 times that of PSO and 9.30 times that of DE. In the single-peak function used above, the average speed of OIO reaches 1.89 times that of PSO and 8.63 times that of DE, and in the multimodal function, the average speed of OIO reaches 1.59 times that of PSO and 9.63 times that of DE. Therefore, we can infer that OIO outperforms DE and PSO in terms of speed, whether it is used to search for multimodal or single-peak functions, and when the threshold is set to 1e-14, the average speed of OIO opens the largest gap with DE, reaching 10.39 times that of DE. And when the threshold is set to 1e-16, the average speed of OIO opens the biggest gap with PSO, reaching 2.27 times of PSO.
6.3 CPU Usage Test Results
Considering the significantly longer time consumption of the DE algorithm compared to PSO and OIO in reaching the acceptable threshold, the comparison of CPU usage rates will no longer include the DE algorithm. To rigorously investigate the impact of algorithms on CPU usage, we first run the threads responsible for recording CPU usage separately 10 times, each for 2 minutes, to obtain the baseline system load. Through measurements from the 10 experiments, the average baseline load is calculated to be 0.65
Table 12: Average Resource Usage
OIO PSO
Test1 4.340299 4.801111
Test2 4.332787 4.686250
Test3 4.616883 4.780220
Test4 4.806944 4.761538
Test5 4.460274 4.845455
Test6 4.464634 4.739560
Test7 4.746667 4.690217
Test8 4.496667 4.628889
Test9 4.271698 4.412088
Test10 4.501124 4.722093
Table 13: Resource Consumption After Debasement
OIO PSO
Test1 3.690299 4.151111
Test2 3.682787 4.036250
Test3 3.966883 4.130220
Test4 4.156944 4.111538
Test5 3.810274 4.195455
Test6 3.814634 4.089560
Test7 4.096667 4.040217
Test8 3.846667 3.978889
Test9 3.621698 3.762088
Test10 3.851124 4.072093
In the ten experiments, the average net CPU usage rates for OIO and PSO were ${3.85}\%$ and ${4.05}\%$ , respectively. OIO demonstrates an advantage in CPU usage, achieving faster results without significantly increasing CPU computational resources. This efficiency arises from OIO’s optimized scheduling and resource allocation, enhancing computational efficiency and reducing CPU load. Consequently, even during high-intensity computing tasks, the system can maintain lower resource consumption. This characteristic makes OIO more competitive in scenarios requiring rapid processing of large amounts of data with limited resources. Therefore, OIO not only outperforms PSO in computational speed but also optimizes CPU resource usage, making it suitable for scenarios with limited computational resources or requiring efficient energy consumption ratios.
Figure 11: CPU total usage area chart
Figure 11 presents a comparison of the total CPU usage between OIO and PSO during the first experiment. We can visually observe that the peak CPU usage of the OIO algorithm is smaller than that of PSO, and OIO also terminates the search process earlier. The frequent significant spikes in CPU usage observed during the search process of PSO further confirm the superiority of the OIO algorithm in reducing resource usage.
6.4 Engineering problem performance
6.4.1 The Three-member Truss Design Problem
Figure 12: Mean Convergence Curve-Three-Bar Truss Design Problems
Table 10: Indicator Statistics of Problems with Three-Rod Truss Design
Algorithms OIO PSO WOA DE OOA HBA DBO COA
Optimum Value 263.9072 263.9106 263.9187 263.8996 264.2354 263.8985 264.2569 263.925
Worst Value 263.9571 263.9853 274.8401 264.6185 267.4255 264.1518 270.0506 265.0755
Standard Deviation 0.018585 0.027862 4.52559 0.303941 1.205465 0.10063 2.253484 0.473283
Mean Value 263.9293 263.94 267.0542 264.0805 265.4593 264.0287 266.4151 264.2431
Median Value 263.9311 263.9361 264.7892 263.9435 265.0254 263.9966 265.9192 264.0707
In Table 10, the OIO algorithm outperformed all other algorithms in the three-member truss design problem. It achieved the lowest optimal value of 263.9072 and the smallest standard deviation of 0.018585, demonstrating exceptional stability and consistency. Additionally, as illustrated in Figure 12, OIO exhibited the fastest convergence speed among the evaluated algorithms. In contrast, the DBO algorithm ranked last, presenting the slowest convergence and the highest standard deviation of 2.253484, indicating significant variability. OOA and WOA followed as the second and third worst performers, with standard deviations of 1.205465 and 4.52559, respectively. The remaining algorithms—PSO, DE, HBA, and COA—showed similar performance levels with comparable standard deviations and optimal values. These findings clearly establish OIO as the most effective and robust algorithm for the three-member truss design problem, surpassing all other methods in both precision and reliability.
6.4.2 The Tension/Compression Spring Design Problem
Figure 13: Average Convergence Curve - Compression Spring Design
Table 11: Compression Spring Design Index Statistics
Algorithms OIO PSO WOA DE OOA HBA DBO COA
Optimum Value 0.012739 0.013114 0.012682 0.013088 0.012974 0.012728 0.0133 0.013019
Worst Value 0.012882 0.01471 0.015443 0.029825 3.37E+14 0.018059 0.030467 2.38E+13
Standard Deviation 6.06E-05 0.000612 0.001084 0.007244 1.45E+14 0.002647 0.006545 1.06E+13
Mean Value 0.012818 0.013703 0.013639 0.02085 9.05E+13 0.015146 0.024234 4.75E+12
Median Value 0.012803 0.01349 0.013319 0.020402 7.62E+12 0.014173 0.025 0.013481
In Table 11 and Figure 13, the results of the tension/compression spring design problem highlight the significant advantages of the OIO algorithm. While OIO achieved the second-fastest convergence speed, it ultimately obtained the best optimal value (0.012739) among the eight algorithms tested. Additionally, OIO demonstrated exceptional stability and consistency, with a low standard deviation (6.06E-5) and a narrow range between its best and worst values (0.012739 to 0.012882). In comparison, algorithms such as OOA and COA showed significantly poorer performance, with their results deviating substantially from the optimal values. DBO exhibited large errors and fluctuations, while DE converged much slower. Although HBA converged faster than OIO, its optimal value (0.012728) was slightly inferior. These results underscore OIO’s ability to balance convergence speed, precision, and robustness, making it the most reliable and effective algorithm for solving this problem.
7. Conclusion and Discussion
This paper presents a novel bionic intelligence optimisation algorithm, the Octopus Inspired Optimization (OIO) algorithm, inspired by the unique neuromorphic features of the octopus, including hierarchical structure, decentralised interactions, and region regeneration. The design of the OIO algorithm not only reflects a deep understanding of individual and collective intelligence, but also has important foresight in the field of heuristic algorithms.
7.1 Performance and Efficiency
Experimental results show that the OIO algorithm outperforms established algorithms such as Particle Swarm Optimisation (PSO) and Differential Evolution (DE) in terms of speed, accuracy and robustness. In particular, at a tighter threshold 1e-16, OIO achieves an average speed of 2.27 times that of PSO and 9.63 times that of DE on multimodal functions. For example, in the test of the Griewank function, when the threshold is set to 1e-16, the convergence time of the OIO is 4.53 seconds, which is much better than that of the PSO at 7.24 seconds and that of the DE at 29.12 seconds. In addition, when the threshold is set to 1e-14, the average speed of OIO opens the largest gap with DE, reaching 10.39 times that of DE.
7.2 Algorithm Adaptation and Resource Utilisation
A distinguishing feature of the OIO algorithm is its adaptability. The algorithm effectively balances exploration and exploitation through its multi-layer search and asynchronous parallelism mechanisms, adapting to a wide range of optimisation problems. In addition, the region regeneration mechanism enhances the algorithm’s ability to escape from local optima by allowing less adaptable tentacles to randomly regenerate, thus promoting diversity in the search process.
OIO demonstrates significant resource efficiency in CPU usage tests. In ten experiments, OIO has a lower average CPU usage compared to PSO, 3.85% versus 4.05%, respectively. This result shows that OIO does not additionally use more CPU resources while getting results faster, showing its energy saving and efficiency.
7.3 Directions for Follow-up Research
The success of OIO raises several implications for the field of optimisation. Firstly, it highlights the potential for bio-inspired algorithms to be more effective than traditional methods in solving complex and dynamic problems. Second, the principles demonstrated by OIO can be extended to other domains, such as robotics and decision support systems, where adaptive and efficient problem solving capabilities are valuable.
For future research, exploring the combination of OIO with machine learning approaches may enhance its predictive capabilities, allowing for more efficient handling of dynamically changing environments. In addition, extending the algorithm to handle multi-objective optimisation problems may significantly expand its application in various disciplines where navigating trade-offs between conflicting objectives is required.
7.4 Conclusion
In summary, the Octopus Inspired Optimization (OIO) algorithm represents an important advancement in the field of heuristic algorithms, providing a robust, efficient, and adaptable approach to solving optimisation problems. Its design is inspired by the complex neuromorphic structure of the octopus, providing a new perspective on the synthesis of individual and collective intelligence in artificial systems. The continued development and application of OIO is expected to have a far-reaching impact in a number of areas of theoretical research and practical problem solving.
Appendix
Code Open-source Statement
In our research, we have utilized and developed the Octopus Inspired Optimization (OIO) algorithm to address multi-objective optimization problems. To promote academic exchange and further research, we have open-sourced this code on GitHub. Readers and researchers can access and download the code through the link: https://github.com/JLU-WangXu/Octopus_Inspired_Optimization_OIO
Table 4: OIO elapsed time (time/s)
Function Acceptable 1.00E-08 1.00E-10 1.00E-12 1.00E-14 1.00E-16
Griewank Mean 5.741724 6.188067 6.36652 6.999679 4.533972
Std./Mean 0.812342 0.912028 0.714705 0.745545 0.534402
Levy Mean 1.684242 1.400081 1.398468 1.429915 1.806029
Std./Mean 0.567357 0.125137 0.019156 0.009957 0.499387
Rastrigin Mean 1.10316 1.194469 1.135208 1.113722 1.657665
Std./Mean 0.150621 0.034087 0.192189 0.193329 0.637898
Dixon_Price Mean 1.406462 0.963744 1.453894 1.044185 1.074503
Std./Mean 0.651836 0.150012 0.622538 0.014187 0.01571
Sum_Square Mean 0.737754 0.791872 0.813084 0.832736 0.898983
Std./Mean 0.303963 0.267169 0.24071 0.260441 0.199391
Trid Mean 1.735214 1.464731 2.011118 1.964295 1.581761
Std./Mean 0.427796 0.107016 0.489038 0.509212 0.013935
Rosenbrock Mean 1.007514 1.66239 4.230964 8.0003 9.78229
Std./Mean 0.668706 0.545263 0.645342 0.397721 0.057744
Sphere Mean 0.939359 0.966379 0.946484 0.951027 0.985048
Std./Mean 0.00851 0.009343 0.138287 0.137066 0.137804
Rotated_Hyper_Ellipsoid Mean 0.811142 0.870507 0.804912 0.904883 0.852174
Std./Mean 0.220457 0.154054 0.258098 0.142306 0.239823
Sum_Of_Different_Powers Mean 0.931518 1.031799 1.082279 1.071653 1.123941
Std./Mean 0.140066 0.014347 0.017394 0.123393 0.11667
Schaffer_Function_N.2 Mean 1.01016 1.057394 1.056161 1.05103 1.071514
Std./Mean 0.021466 0.010542 0.023057 0.141186 0.128248
Beale Mean 5.570163 5.487116 4.170626 5.267392 3.904259
Std./Mean 0.59432 0.493344 0.792101 0.746996 0.887484
Drop_Wave Mean 0.96957 1.06181 1.38016 1.325116 1.204039
Std./Mean 0.193394 0.119166 0.525539 0.569184 0.058057
Levy_N.13 Mean 0.963707 1.509239 1.74332 1.088987 1.290307
Std./Mean 0.15146 0.634362 0.624015 0.017901 0.564722
Easom Mean 1.080786 1.357562 1.365644 1.112898 1.143276
Std./Mean 0.015286 0.560356 0.538549 0.137081 0.130198
Table 5: Error table for changing the number of suction cups
Function Metric 30,200 40,200 50,200 60,200 70,200
Ackley Mean 4.44E-16 4.44E-16 4.44E-16 4.44E-16 4.44E-16
Std. 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
Griewank Mean 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
Std. 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
Levy Mean 1.50E-32 1.50E-32 1.50E-32 1.50E-32 1.50E-32
Std. 2.88E-48 2.88E-48 2.88E-48 2.88E-48 2.88E-48
Rastrigin Mean 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
Std. 0 0 0 0 0
Schwefel Mean -2.30E-13 -2.30E-13 -2.30E-13 -2.30E-13 -2.30E-13
Std. 0 0 0 0 0
Dixon_Price Mean 2.47E-32 2.47E-32 2.47E-32 2.47E-32 2.47E-32
Std. 5.77E-48 5.77E-48 5.77E-48 5.77E-48 5.77E-48
Sum_Square Mean 1.20E-76 1.03E-78 1.21E-77 1.71E-78 2.26E-76
Std. 2.16E-76 1.11E-78 1.38E-77 4.37E-78 5.50E-76
Trid Mean 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
Std. 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
Rosenbrock Mean 7.70E-31 7.96E-31 1.40E-30 9.68E-31 1.21E-30
Std. 8.44E-31 2.08E-30 2.79E-30 1.64E-30 2.28E-30
Sphere Mean 3.76E-62 1.83E-61 4.55E-62 1.13E-61 1.54E-61
Std. 3.98E-62 2.20E-61 2.96E-62 1.20E-61 1.31E-61
Rotated_Hyper_Ellipsoid Mean 1.49E-76 5.69E-76 5.28E-74 3.50E-76 3.20E-76
Std. 2.05E-76 1.17E-75 1.09E-73 4.21E-76 6.45E-76
Sum_Of_Different_Powers Mean 5.77E-38 3.79E-38 4.04E-38 5.04E-38 1.62E-37
Std. 1.14E-37 5.17E-38 7.62E-38 1.06E-37 2.46E-37
Schaffer_Function_N.2 Mean 0 0 0 0 0
Std. 0 0 0 0 0
Z Beale Mean 0 0 0 0 0
Std. 0 0 0 0 0
Goldstein_Price Mean -8.10E-14 -8.10E-14 8.05E-14 -8.00E-14 -8.00E-14
Std. 1.93E-16 2.11E-16 1.69E-16 7.78E-16 1.26E-16
Drop_Wave Mean 0 0 0 0 0
Std. 0 0 0 0 0
Levy_N.13 Mean 1.35E-31 1.35E-31 1.35E-31 1.35E-31 1.35E-31
Std. 2.31E-47 2.31E-47 2.31E-47 2.31E-47 2.31E-47
Easom Mean 0 0 0 0 0
Std. 0 0 0 0 0
Table 6: Error table for changing the number of iterations
Function Metric 50,160 50,180 50,200 50,220 50,240
Ackley Mean 4.44E-16 4.44E-16 4.44E-16 4.44E-16 4.44E-16
Std. 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
Griewank Mean 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
Std. 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
Levy Mean 1.50E-32 1.50E-32 1.50E-32 1.50E-32 1.50E-32
Std. 2.88E-48 2.88E-48 2.88E-48 2.88E-48 2.88E-48
Rastrigin Mean 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
Std. 0 0 0 0 0
Schwefel Mean -2.3E-13 -2.3E-13 -2.3E-13 -2.3E-13 -2.3E-13
Std. 0 0 0 0 0
Dixon_Price Mean 2.47E-32 2.47E-32 2.47E-32 2.47E-32 2.47E-32
Std. 5.77E-48 5.77E-48 5.77E-48 5.77E-48 5.77E-48
Sum_Square Mean 4.72E-78 4.21E-77 1.21E-77 1.04E-76 7.64E-77
Std. 7.78E-78 1.05E-76 1.38E-77 2.08E-76 1.60E-76
Trid Mean 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
Std. 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
Rosenbrock Mean 9.32E-31 8.01E-31 1.40E-30 1.84E-31 5.69E-31
Std. 1.48E-30 9.35E-31 2.79E-30 2.78E-31 7.85E-31
Sphere Mean 2.05E-61 1.02E-61 4.55E-62 6.51E-62 1.04E-61
Std. 4.20E-61 1.09E-61 2.96E-62 8.57E-62 1.31E-61
Rotated_Hyper_Ellipsoid Mean 2.37E-76 4.57E-76 5.28E-74 4.18E-76 3.67E-76
Std. 4.85E-76 6.27E-76 1.09E-73 5.23E-76 1.12E-75
Sum_Of_Different_Powers Mean 1.59E-37 8.35E-38 4.04E-38 4.94E-38 1.50E-37
Std. 3.82E-37 1.36E-37 7.62E-38 8.02E-38 3.02E-37
Schaffer_Function_N.2 Mean 0 0 0 0 0
Std. 0 0 0 0 0
Beale Mean 0 0 0 0 0
Std. 0 0 0 0 0
Goldstein_Price Mean -8E-14 -8.1E-14 8.05E-14 -8E-14 -8E-14
Std. 1.69E-16 2.11E-16 1.69E-16 3.54E-16 3.95E-16
Drop_Wave Mean 0 0 0 0 0
Std. 0 0 0 0 0
Levy_N.13 Mean 1.35E-31 1.35E-31 1.35E-31 1.35E-31 1.35E-31
Std. 2.31E-47 2.31E-47 2.31E-47 2.31E-47 2.31E-47
Easom Mean 0 0 0 0 0
Std. 0 0 0 0 0
Table 7: Table of elapsed time (Acceptable: 1e-8)
Function Metric PSO DE OIO AIS GA ABC
Griewank Mean 4.150111 24.76257 5.741724 22.71635 308.1214 650.3767
Std./Mean 0.129429 0.023299 0.812342 0.001335 0.000253 0.001058
Levy Mean 1.353939 6.764352 1.684242 12.94923 197.457 510.7123
Std./Mean 0.195083 0.072718 0.567357 0.047874 0.016349 0.001055
Rastrigin Mean 1.682712 11.09525 1.10316 6.611436 141.8939 424.2892
Std./Mean 0.101307 0.043616 0.150621 0.002962 0.013819 0.001309
Dixon_Price Mean 1.161798 11.72637 1.406462 2.723215 101.751 362.4495
Std./Mean 0.111377 0.109422 0.651836 0.006024 0.018707 0.001459
Sum_Square Mean 0.294829 1.859835 0.737754 0.082874 58.79886 2.189965
Std./Mean 0.442125 0.265529 0.303963 1.097617 0.733724 0.311005
Trid Mean 1.032021 5.812206 1.735214 3.948256 113.3602 397.34
Std./Mean 0.108503 0.092045 0.427796 0.003102 0.000537 0.003336
Rosenbrock Mean 1.813837 14.37608 1.007514 2.741521 100.7789 358.5214
Std./Mean 0.091199 0.042316 0.668706 0.007522 0.000259 0.001759
Sphere Mean 0.937772 5.22159 0.939359 2.643382 100.3239 355.6785
Std./Mean 0.05752 0.08406 0.00851 0.013105 0.009964 0.001872
Rotated_Hyper_Ellipsoid Mean 0.487025 2.459416 0.811142 0.100037 79.33034 3.367999
Std./Mean 0.288342 0.302576 0.220457 0.767783 0.370295 0.329116
Sum_Of_Different_Powers Mean 1.474839 7.293044 0.931518 1.851535 91.07604 341.1465
Std./Mean 0.135683 0.063349 0.140066 0.01159 0.023916 0.001331
Schaffer_function_N.2 Mean 1.040904 17.69117 1.01016 4.52061 119.0329 380.94
Std./Mean 0.106365 0.071751 0.021466 0.002505 0.000314 0.001219
Beale Mean 1.090452 8.004226 5.570163 3.407871 107.4548 351.8095
Std./Mean 0.084624 0.077231 0.59432 0.004779 0.031899 0.002247
Drop_Wave Mean 1.505673 20.33826 0.96957 5.640463 130.9503 387.2101
Std./Mean 0.066389 0.02885 0.193394 0.003744 0.037512 0.001539
Levy_N.13 Mean 1.282261 7.457867 0.963707 3.771574 111.3559 372.5849
Std./Mean 0.094228 0.021857 0.15146 0.001279 0.023494 0.001603
Easom Mean 1.613746 14.00182 1.080786 6.323894 138.4347 406.2244
Std./Mean 0.146929 0.033615 0.015286 0.001705 0.006713 0.001152
Table 8: Table of elapsed time (Acceptable: 1e-10)
Function Metric PSO DE OIO AIS GA ABC
Griewank Mean 5.120116 26.6586 6.188067 23.75453 326.3725 750.3454
Std./Mean 0.165551 0.026642 0.912028 0.042972 0.029374 0.00092
Levy Mean 2.156131 8.630126 1.400081 13.35455 210.41 593.0117
Std./Mean 0.071349 0.057354 0.125137 0.04991 0.022777 0.000933
Rastrigin Mean 2.164205 13.23251 1.194469 6.820684 149.4181 495.4236
Std./Mean 0.067422 0.02743 0.034087 0.001938 0.017215 0.001214
Dixon_Price Mean 1.560746 14.82201 0.963744 2.693343 104.5691 424.1581
Std./Mean 0.142253 0.093025 0.150012 0.003187 0.020812 0.001901
Sum_Square Mean 0.487518 3.067631 0.791872 0.825671 95.75756 4.2847
Std./Mean 0.456657 0.110409 0.267169 1.107406 0.294717 0.476724
Trid Mean 1.44406 7.604722 1.464731 4.136783 120.2932 460.4977
Std./Mean 0.086068 0.104111 0.107016 0.00104 0.027689 0.001312
Rosenbrock Mean 2.427417 17.32397 1.66239 2.733574 104.3866 422.1888
Std./Mean 0.105989 0.023064 0.545263 0.021404 0.014805 0.001283
Sphere Mean 1.332293 7.076579 0.966379 2.777539 105.5117 420.2953
Std./Mean 0.072704 0.064156 0.009343 0.033969 0.014218 0.000776
Rotated_Hyper_Ellipsoid Mean 0.769272 3.574589 0.870507 1.058892 96.62095 9.481768
Std./Mean 0.291967 0.155919 0.154054 0.826095 0.013199 0.183518
Sum_Of_Different_Powers Mean 2.23992 9.985262 1.031799 1.790721 92.69818 407.6053
Std./Mean 0.073515 0.022813 0.014347 0.001579 0.010642 0.001089
Schaffer_function_N.2 Mean 1.472666 20.84638 1.057394 4.478356 125.6203 451.7385
Std./Mean 0.062299 0.020838 0.010542 0.003539 0.013461 0.000684
Beale Mean 1.538096 11.03384 5.487116 3.443204 110.9274 411.3184
Std./Mean 0.09564 0.033855 0.493344 0.02687 0.011191 0.001504
Drop_Wave Mean 1.890358 21.70964 1.06181 5.536673 136.0943 412.6167
Std./Mean 0.111611 0.054839 0.119166 0.001858 0.021377 0.078579
Levy_N.13 Mean 1.716771 9.159263 1.509239 3.643332 116.0413 375.4466
Std./Mean 0.097177 0.078981 0.634362 0.002394 0.014216 0.001116
Easom Mean 2.196847 15.77514 1.357562 6.203204 142.2855 412.1029
Std./Mean 0.064935 0.043593 0.560356 0.002803 0.010149 0.000392
Table 9: Table of elapsed time (Acceptable: 1e-12)
Function Metric PSO DE OIO AIS GA ABC
Griewank Mean 5.758299 29.04583 6.36652 22.71843 320.4427 642.5988
Std./Mean 0.093315 0.038646 0.714705 0.001661 0.003166 0.000707
Levy Mean 2.724082 9.928845 1.398468 12.34368 201.4017 506.4917
Std./Mean 0.068631 0.071212 0.019156 0.002691 0.002483 0.001381
Rastrigin Mean 2.741055 14.85306 1.135208 6.518731 144.5299 420.5592
Std./Mean 0.055742 0.04426 0.192189 0.001236 0.005335 0.001339
Dixon_Price Mean 2.240487 19.48927 1.453894 2.724012 103.3059 359.5647
Std./Mean 0.068441 0.097374 0.622538 0.00165 0.00312 0.001023
Sum_Square Mean 0.842899 3.897546 0.813084 2.739547 102.8596 8.685232
Std./Mean 0.375862 0.209218 0.24071 0.001655 0.004469 0.247585
Trid Mean 1.92432 9.877959 2.011118 3.923895 118.6116 393.2193
Std./Mean 0.069481 0.0843 0.489038 0.001014 0.005168 0.001561
Rosenbrock Mean 3.461905 20.70815 4.230964 2.74874 102.5651 359.026
Std./Mean 0.077531 0.030162 0.645342 0.001861 0.003454 0.000765
Sphere Mean 1.812573 8.656818 0.946484 2.602898 101.9811 356.5148
Std./Mean 0.053603 0.043488 0.138287 0.003237 0.005663 0.001663
Rotated_Hyper_Ellipsoid Mean 1.1708 5.045653 0.804912 2.019528 93.1532 11.07505
Std./Mean 0.211255 0.062912 0.258098 0.004327 0.003877 0.206102
Sum_Of_Different_Powers Mean 3.087205 13.00794 1.082279 1.831578 90.79551 340.2233
Std./Mean 0.051939 0.018289 0.017394 0.005751 0.003216 0.001151
Schaffer_function_N.2 Mean 1.889367 22.65672 1.056161 4.433171 119.6673 381.5241
Std./Mean 0.071266 0.034615 0.023057 0.001642 0.0038 0.000754
Beale Mean 2.065522 13.40765 4.170626 3.41407 107.3414 354.7624
Std./Mean 0.045787 0.059634 0.792101 0.002669 0.00293 0.001151
Drop_Wave Mean 2.51205 23.96628 1.38016 5.541868 128.8584 388.5166
Std./Mean 0.083472 0.030999 0.525539 0.002167 0.004874 0.001792
Levy_N.13 Mean 2.370653 10.96674 1.74332 3.785729 110.9691 373.6242
Std./Mean 0.047936 0.093355 0.624015 0.00167 0.003777 0.000731
Easom Mean 2.661788 17.96128 1.365644 6.320979 137.3055 407.3709
Std./Mean 0.089779 0.025593 0.538549 0.004496 0.002609 0.000656
Table 10: Table of elapsed time (Acceptable: 1e-14)
Function Metric PSO DE OIO AIS GA ABC
Griewank Mean 6.501565 31.83202 6.999679 23.45174 321.5481 640.1049
Std./Mean 0.067926 0.018931 0.745545 0.001231 0.028744 0.001482
Levy Mean 3.277244 12.85423 1.429915 12.80924 208.8177 507.7562
Std./Mean 0.071585 0.024544 0.009957 0.002518 0.024776 0.001425
Rastrigin Mean 3.111076 17.11876 1.113722 6.880304 147.0575 421.9303
Std./Mean 0.050817 0.035539 0.193329 0.00442 0.001316 0.001066
Dixon_Price Mean 2.447319 21.73384 1.044185 2.823881 106.298 360.3045
Std./Mean 0.069938 0.122666 0.014187 0.003067 0.016454 0.000596
Sum_Square Mean 1.199011 5.019165 0.832736 2.840054 105.379 13.22378
Std./Mean 0.16827 0.091475 0.260441 0.001713 0.017287 0.218352
Trid Mean 2.375264 12.13656 1.964295 4.140257 119.7874 393.7792
Std./Mean 0.036267 0.044306 0.509212 0.002638 0.021225 0.00125
Rosenbrock Mean 4.103158 23.87009 8.0003 2.813445 105.8452 359.1283
Std./Mean 0.047667 0.050567 0.397721 0.00193 0.019351 0.000448
Sphere Mean 2.126858 10.72363 0.951027 2.674969 104.8224 357.4763
Std./Mean 0.077852 0.040592 0.137066 0.002481 0.024217 0.001172
Rotated_Hyper_Ellipsoid Mean 1.535709 5.68187 0.904883 2.078267 97.87827 16.62687
Std./Mean 0.178987 0.125927 0.142306 0.002699 0.017016 0.136557
Sum_Of_Different_Powers Mean 3.644701 15.53834 1.071653 1.853026 93.34616 342.0732
Std./Mean 0.045579 0.018577 0.123393 0.002464 0.018168 0.000843
Schaffer_Function_N.2 Mean 2.444901 24.91335 1.05103 4.587549 123.7221 381.9203
Std./Mean 0.051015 0.027116 0.141186 0.00142 0.00459 0.002107
Beale Mean 2.487051 16.70951 5.267392 3.432029 109.8748 354.1861
Std./Mean 0.048967 0.045322 0.746996 0.001399 0.011214 0.001359
Drop_Wave Mean 2.859051 26.29869 1.325116 5.690425 133.5877 389.2991
Std./Mean 0.059839 0.028364 0.569184 0.001781 0.006785 0.001408
Levy_N.13 Mean 2.665374 12.94997 1.088987 3.974641 111.3743 375.7356
Std./Mean 0.064717 0.032108 0.017901 0.002366 0.001232 0.000879
Easom Mean 3.129896 20.24751 1.112898 6.357879 144.0042 414.082
Std./Mean 0.048175 0.020506 0.137081 0.003913 0.014414 0.001883
Table 11: Table of elapsed time (Acceptable: 1e-16)
Function Metric PSO DE OIO AIS GA ABC
Griewank Mean 7.24397359 29.11755 4.533972 23.84387412 314.5254 666.8623
Std./Mean 0.02502536 0.043305 0.534402 0.002918369 0.033835 0.001468
Levy Mean 3.88878439 10.32649 1.806029 12.3126343 199.5628 521.6451
Std./Mean 0.03818811 0.066476 0.499387 0.001963862 0.000436 0.001547
Rastrigin Mean 3.29537838 15.09525 1.657665 6.677716303 145.2885 434.172
Std./Mean 0.04389171 0.029072 0.637898 0.00498335 0.000436 0.002237
Dixon_Price Mean 2.96643715 18.82834 1.074503 2.698420405 101.5288 377.2782
Std./Mean 0.05310544 0.065887 0.01571 0.002842285 0.000819 0.002939
Sum_Square Mean 1.78254576 4.123675 0.898983 2.727951646 102.0037 9.649681
Std./Mean 0.13507084 0.106175 0.199391 0.011082873 0.000562 0.186023
Trid Mean 2.74078002 9.796526 1.581761 4.017462444 115.4904 409.2938
Std./Mean 0.0482944 0.071439 0.013935 0.001894483 0.00041 0.002042
Rosenbrock Mean 5.15432875 20.24328 9.78229 2.751228285 101.7115 375.8151
Std./Mean 0.10554766 0.033836 0.057744 0.013401405 0.000852 0.002312
Sphere Mean 2.55244815 8.565935 0.985048 2.625390673 101.1165 368.97
Std./Mean 0.04708961 0.056631 0.137804 0.018479112 0.000756 0.002475
Rotated_Hyper_Ellipsoid Mean 1.864943 4.913963 0.852174 1.992052841 94.95243 12.46418
Std./Mean 0.10278724 0.151184 0.239823 0.022651842 0.035155 0.278332
Sum_Of_Different_Powers Mean 4.37806702 12.77827 1.123941 1.758338785 89.80916 357.1348
Std./Mean 0.05397275 0.029037 0.11667 0.004191448 0.013738 0.001533
Schaffer_function_N.2 Mean 2.81845226 22.91213 1.071514 4.489289975 121.939 397.1624
Std./Mean 0.06273317 0.034797 0.128248 0.005297317 0.000508 0.001909
Beale Mean 3.02192376 13.76873 3.904259 3.358512115 108.0523 368.7202
Std./Mean 0.0703482 0.042239 0.887484 0.001552766 0.00065 0.001865
Drop_Wave Mean 3.37993839 24.14356 1.204039 5.567820978 130.7706 402.7617
Std./Mean 0.04742879 0.04462 0.058057 0.005100557 0.000303 0.00191
Levy_N.13 Mean 3.07376905 11.03845 1.290307 3.734168887 110.9802 392.4343
Std./Mean 0.05146962 0.032883 0.564722 0.007472344 0.00069 0.002444
Easom Mean 3.67135949 17.82801 1.143276 6.344906807 139.5839 422.5718
Std./Mean 0.03994282 0.027393 0.130198 0.004826897 0.012542 0.004799